abc is a triangle in which ab = 12 cm , ca = 13cm , c b = 5 cm find m (< abc )
Answer by kietra(57)

ab = 12 cm , ca = 13cm , c b = 5 cm To find m (< abc, you must use the cosine rule. You can use the cosine rule when you are given two sides and the included angle or when you are given three sides.
cosB = (a^2 + c^2 - b^2)/2ac
cosB = (5^2 + 12^2 - 13^2)/2*5*12
cosB = (25 + 144 - 169)/120
cosB = (169 - 169)/120
cosB = 0/120 = 0
B = cos^-1(0)
B = 90 degrees
is there any other help we tutors can help with?